OpenStudy (anonymous):

(2 x 105th power)(2 x 106th power)

OpenStudy (anonymous):

\[(2 *10^{5})2^{106}\] that?

OpenStudy (anonymous):

10 to the 6th power

OpenStudy (anonymous):

so (2*10^5)(2*10^6)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay to do this one is relatively easy and I even double checked it with my calculator. All we need to do is rearrange it like so 2*2*10^5*10^6 that should make it easier for you to see if you still need help let me know but I try to leave as much up to you as possible.

OpenStudy (anonymous):

10^5 = 100000 10^6 = 1000000 multiply them both by 2 200000 * 2000000 = 400000000000

OpenStudy (anonymous):

or add the powers and multiply the others to simply get 4 * 10^11

OpenStudy (anonymous):

yes ^^ that is a much nicer solution

OpenStudy (anonymous):

thanks